OpenStudy (anonymous):

Prove, using the epsilon, delta definition of a limit. Limit as x->3 of 1/(x+3)^4 = positive infinity

OpenStudy (anonymous):

Woops, made a mistake, it should be as x approaches -3. \[\lim_{x \rightarrow -3}\frac{ 1 }{ (x+3)^{4} }=\infty\]